OpenStudy (wwe123):

the roots of the quadratic equation 4x^2-(5a+1)x+5a=0 are p & q. if q=1+p,calculate the possible values of a,p& q.

OpenStudy (ajprincess):

 Similarly we can find a relation for ur eqn. \[4x^2-(5a+1)x+5a=0\] \[4(x^2-\frac{(5a+1)}{4}x+\frac{5a}{4})=0\] \[p+q=\frac{(5a+1)}{4}\] \[p+1+p=\frac{(5a+1)}{4}\] \[2p+1=\frac{(5a+1)}{4}-(1)\] \[4(2p+1)=5a+1\] \[pq=\frac{5a}{4}\] \[p(1+p)=\frac{5a}{4}\] \[4p(1+p)=5a-(2)\] Using equations (1) and (2) find p and a. Can u do it @wwe123?